OpenStudy (anonymous):

Hi guys, I need help solving my last 2 homework problems. The first one is: INT (dy)/ Squareroot(16 + y^2)?

OpenStudy (anonymous):

trig sub or hyperbolic trig function? i bet there is a formula on the last page of your text for this, or rather for \[\int\frac{dx}{\sqrt{a^2+x^2}}\]